OpenStudy (anonymous):

point slope: containing (-2,6) and having slope -3

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

my paper says to give the standard form , graphing form , and point slope idk i guess

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] \[m=-3, x_1=-2,y_1=6\] \[y-6=-3(x+2)\] \[y-6=-3x-6\] \[y=-3x\]

OpenStudy (anonymous):

the one i wrote is slope-intercept form, maybe "graphing form"

OpenStudy (anonymous):

standard form of a line is Ax+By=C point-slope form of a line is y-y1=m(x-x1) graphing form is probably slope-intercept form, which is y=mx+b

OpenStudy (anonymous):

point slope form is \[y-6=-3(x+2)\]

OpenStudy (anonymous):

standard form is \[3x+y=0\]

OpenStudy (anonymous):

y=-3x is the equation and graph for this equation is

OpenStudy (anonymous):

this one

OpenStudy (anonymous):

i just need the point slope now like how do i plot the numbers given?

OpenStudy (anonymous):

y=-3x + 0 is in point slope form with slope is -3 and y intercept is 0

OpenStudy (anonymous):

how you get that?